# The Lunar Gravitational-wave Antenna: Mission Studies and Science Case

### v1

@@ 10 Mpc        30 Mpc        100 Mpc
Several studies have assessed the merger rates of MBH binaries as a function of the mass of the merging bodies and redshift [509, 510] employing semi-analytical models for the cosmic evolution of galaxies and MBHs. The obtained predictions can vary substantially depending on the selected black hole seed mass (light versus heavy seeds) and physical prescriptions implemented to describe the physics of the inspiral; the intrinsic number of MBH mergers is typically of order a few tens to a few hundreds per year, with light-seed models predicting a larger number of events dominated by mergers of MBHs with $<10^{4}M_{\odot}$, while the heavy seed models tend to feature fewer events, mainly involving MBHs of $10^{4}-10^{6}M_{\odot}$. The LISA GW detector (which is sensitive to larger masses and higher redshift, if compared with LGWA) will likely have its detection rates (again of tens to hundreds of events per year) dominated by mergers occurring at $z\sim 5$ [510]; this suggests that the LGWA will be able to detect at least a few (tens) of events, and possibly have its detection rates dominated by lower-$z$ mergers bridging the gap.
@@
LGWA is going to predominately observe TDEs occurring about relatively low-mass MBHs ($<10^{6}M_{\odot}$), likely inhabiting dwarf galaxies or globular clusters in the local Universe [489]. This poses a series of challenges for the proper modelling of the rates, as the occupation fraction of MBHs in those low-mass systems remains debated [568], as well as their innermost stellar density profile (which remains virtually unresolvable at the scale of a fraction of the MBH influence radius [569], where most of TDEs are expected to be generated [570]). Although earlier works have suggested TDE rates would be larger about low mass MBHs [562, 571], this idea has been recently revised and it seems instead that rates about low-mass MBHs drop significantly over time after an initial burst [570, 572]. Overall, the event rates about low-mass MBHs in the local universe remain under close investigation, and upcoming surveys as LSST/Vera Rubin Observatory as well as ULTRASAT and eROSITA will greatly improve our understanding of this in the next few years, thus allowing to refine the science we expect to obtain from the LGWA.
@@ 3.2.3 GW cosmology
where $\Omega_{m,0}$ and $\Omega_{r,0}$ denote the present-day values of the density fractions for matter and radiation, respectively, while $\rho_{\rm DE}(z)$ is the dark energy density as a function of redshift and $\rho_{0}=3H_{0}^{2}/(8\pi G)$. The crucial aspect in the GW context is then obtaining information on the sources’ redshift. In the best-case scenario, the GW event is followed by a detectable EM counterpart which leads to a precise estimate of $z$, as we will discuss in the following. In case a counterpart is not available, GW detections can still be used for cosmological purposes through statistical methods. As first proposed in [582], it is possible to statistically obtain redshift measurements from potential host galaxies falling within the observed GW’s localization error volumes exploiting galaxy catalogs. Also, a reconstruction of the intrinsic source-frame mass distribution of the observed mergers, in particular in the presence of sharp features, can lead to a statistical determination of the sources’ redshifts, being the observed quantities in the detector-frame redshifted. Several recent studies have successfully tested these methods and their combination, as well as other similar ones, with both real data [592, 596, 606, 593, 595, 594, 597, 612] and simulations [587, 589, 618, 619, 620, 621, 622, 623, 624]. For the former method to be informative, detections with narrow error volumes, both in the angular and distance components, are extremely valuable. Another independent approach to measure the expansion history of the Universe as well as the GW bias parameter is using the cross-correlation technique, which can measure the spatial clustering of GW sources with the galaxies [625, 599, 590, 601, 626, 627, 628, 629, 605, 630]. This method can further explore weak gravitational lensing [600, 631, 632] and provide tests of GR modifications at cosmological scales in a model-independent way [627]. Other methods to employ GW detections for cosmological purposes include, e.g., exploiting multiple images of strongly lensed events [633, 634, 635]; having prior knowledge on the redshift distribution of the observed sources [636, 637]; using tidal distortions of NSs combined with EoS knowledge to extract information on the source-frame masses [586, 638, 639]; or measuring the angular barion acoustic oscillation scale from a high number of well localized events [640].
@@
For GW-only events, the so-called ‘dark sirens’ without an EM counterpart and the host galaxy identification, statistical approaches can be adopted. Despite the low statistics, it has also been possible to get an interesting accuracy on the local expansion rate of the Universe [592, 593, 594, 595, 596, 597], as well as other modified gravity observables [596, 610, 614, 611]. As an example, from the binary black-hole merger GW170814 in combination with the Dark Energy Survey it has been possible to set the constraint $H_{0}=75\begin{subarray}{c}+40\\ -32\end{subarray}~{}{\rm km\,s}^{-1}\,{\rm Mpc}^{-1}$ [593].
@@
In particular, in the bright siren case, we expect EM counterparts from CBCs in the Hz-kHz band to be produced by mergers involving at least one deformable object, such as a neutron star, thus coalescences at $\gtrsim{\cal O}({\rm kHz})$ frequencies, falling outside the LGWA band. However, LGWA is able to detect the early inspiral of such systems and can be crucial to localize the source in the sky with good accuracy, hence improving the chances of counterpart identification (see Sec. 2.2.4). Even though the merit of joint GW+EM detections for cosmology is particularly high, systems with an associated counterpart are expected to be rare, with ${\cal O}({\rm few})$ forecasted BNS multimessenger detections per year at LGWA (see Sec. 2.2.1).
@@
In the dark siren case instead, at least for some of the coalescences falling in the band of terrestrial detectors, in particular at low redshifts, LGWA can play a crucial role, significantly shrinking the sky localisation area thanks to the long time spent by the sources in the dHz detector band, and thus the huge effective baseline (see Sec. 2.2.4). Combined with the accuracy in particular third generation terrestrial interferometers can have in the reconstruction of the luminosity distance, such detections can be a game changer for dark-siren GW cosmology, reducing the number of potential host galaxies in such small error volumes to only a few, and thus acting effectively almost as bright sirens. Moreover, the detections of MBH and IMBH mergers that occur in the LGWA band could be used for both the aforementioned methodologies, increasing the statistics and thus the constraining power, with the statistical uncertainty scaling as $\propto 1/\sqrt{N_{\rm events}}$. Another interesting opportunity offered by these massive sources is the possible detection of EM radiation preceding or accompanying the gravitational signal from the environment surrounding the compact objects, making them bright siren candidates highly valuable for cosmology (see Sec. 3.3.4). Finally, as shown in Fig. 7, LGWA will be able to detect coalescences of DWD and NSWD binaries up to $\sim\!100~{}{\rm Mpc}$. Thanks to the long time this kind of sources spend inspiralling in the detector’s band, they can potentially be localized with very high accuracy, and provide a powerful tool to estimate the $H_{0}$ value in the local Universe, both as bright (see Sec. 3.3.1 for a discussion regarding the EM transients these systems can give rise to) and dark sirens in combination with galaxy catalogs if counterpart detections are not available. The estimation of all the other cosmological parameters impacting the $d_{L}-z$ relation at higher redshifts coming from the other classes of sources discussed above would greatly benefit from this low redshift measurement.
@@
Another observable and yet elusive source for current detectors is a SGWB of cosmological origin (CGWB). This refers to tensor modes possibly produced by various physical processes happening in the early universe, including, e.g., phase transitions, inflation, cosmic strings or primordial black holes (see e.g. [644, 645, 646] for reviews). It is, in a sense, the GW equivalent of the Cosmic Microwave Background EM radiation, whose discovery in 1965 has been one of the most relevant in cosmology. The detection of a CGWB would also bring outstanding scientific value, providing direct information about our Universe in its very first instants of evolution that would not be accessible by other means. The characteristic frequency $f_{0}$ of a CGWB at present time is determined by the Hubble rate at the time of generation of the background, redshifted to the present time [645]
@@ 3.3 Multi-messenger observations
The LGWA science case is particularly rich in astrophysical and matter phenomena. As a consequence, many of the GW sources are expected to emit detectable EM signals as well. Most of these events are specific to the deci-Hertz band. Some of these are low-probability events that may be extremely rewarding to detect. In this section, we outline the multi-messenger context of LGWA signals. The prospects of EM observations of binaries containing WDs is described in Sec. 3.3.1. The detection of SNe together with associated GW signals would be a major breakthrough (Sec. 3.3.2). Rare signals observed in EM and GW might come from TDEs of WDs near the horizon of BHs (Sec. 3.3.3). Another multi-messenger study concerns MBH binaries in AGNs, where X-ray observations can complement the GW detection (Sec. 3.3.4). Finally, a new class of EM signals called QPEs are best explained by X-ray emissions from gas in E/IMRI systems (Sec. 3.3.5).
@@ 3.3.1 Compact binaries with white dwarfs and neutron stars
Mergers involving double WD binaries and NSWD binaries can lead to explosive or transient events and the formation of exotic objects that typically do not emerge from the evolutionary path of single stars. Many studies have delved into the mergers of double WD binaries, mostly in the context of potential type Ia supernova progenitors, as reviewed by various authors such as [804, 805, 806, 800, 807]. These binaries are created through common envelope evolution in binary systems, and in the closest double WD binaries, the emission of GWs causes their orbits to shrink and initiates mass transfer between the WDs within a Hubble timescale. The critical factor influencing the final result is whether mass transfer remains stable or diverges, potentially leading to the tidal disruption of the donor star and binary merger [808]. Mass transfer’s stability depends on the response of the WDs to the process, their masses, and various angular momentum transport mechanisms, such as torques due to accretion discs or tidal bulges [809, 810]. Subsequent evolution after WD mergers can yield diverse outcomes, including He-rich hot subdwarfs, R Coronae Borealis stars, massive carbon/oxygen or oxygen/neon WDs, Type Ia supernovae, accretion-induced collapse to a neutron star or even a black hole. A chart summarizing these potential outcomes of WD mergers can be found in Fig. 3 of [811]. The case of type Ia SN holds broader historical significance and particular scientific interest for LGWA, and is therefore discussed in detail in Sec. 3.3.2.
@@
Mergers involving NSWD and BHWD binaries are also expected to fall within the LGWA sensitivity band and generate electromagnetic transients. However, simulating theoretical predictions about their observational signatures to estimate them numerically presents more challenges. The main challenge lies in dealing with a wide range of scales and complex physical processes inherent in this problem. This is in contrast to mergers involving objects of similar size, such as DWD or NSNS/BH, which are more manageable and have more well-developed observational predictions, as demonstrated in studies like [812, 813, 814]. The merger events involving NS or BH with WD have been extensively studied through various hydrodynamical and nuclear-hydrodynamical simulations. Some notable references in this research include [815, 265, 816, 268, 270]. It is generally anticipated that all NSWD mergers result in unstable mass transfer [817]. In such case, the WD is tidally disrupted on dynamical timescales, forming an extended debris disk around the NS. The evolution of this disk is primarily governed by viscosity, but nuclear burning also plays a significant role. Together, disk viscosity and nuclear burning drive outflows throughout the disk [815]. Nuclear burning within the disk progresses steadily, potentially with a weak detonation, and it results in the production of small amounts of 56Ni, typically at most 10-2 M [816]. This limited amount of 56Ni is expected to lead to optical transients, but they are significantly fainter compared to typical type Ia supernovae, which generate over an order of magnitude more 56Ni. The transients resulting from NSWD mergers could potentially represent an entirely distinct class of supernovae. These events might be observable primarily in nearby galaxies using large telescopes or, perhaps, with next-generation survey instruments like Rubin-LSST.
@@
Binaries with a WD component are not loud emitters of gravitational waves, due to their low masses and long orbital period; the strategy for LGWA to detect them is to identify their quasi-monochromatic emission by integrating it over the whole lifespan of the mission, similarly to what is done for continuous wave searches by current gravitational wave detectors. The presence of eccentricity – expected for NSWD binaries if the NS forms after the WD [800] – gives rise to higher-frequency harmonics of the emission [818, 677], which boost overall emitted power as well as placing it closer to the most sensitive band for LGWA. As a summary of what is discussed in Sec. 2.2.1, deci-Hz gravitational wave sources in our own Milky Way are very likely to be observed regardless of their orientation, while sources in nearby, satellite galaxies might be observed in favorable conditions.
[…]
In the event an NS does not collapse into a BH post-merger, Morán-Fraile et al [270] estimates jet velocities similar to those seen in galactic microquasars ($\gtrsim 0.9c$), leading to an event spectrum that initiates with a soft X-ray or UV flash and transitions into lower frequencies, resembling fast blue optical transients (FBOTs, [820]), potentially associated with NSWD mergers. The merger could generate a prolonged, bright radio signal, with the transient’s luminosity potentially exceeding typical kilonovae due to the ejecta’s mass and velocity, rather than r-process yields, which are obscured in kilonovae by lanthanides. If the NS transitions into a BH, this could result in ultrarelativistic jet speeds, leading to long or ultra-long GRB-like transients, marking a significant observational shift

### v2

@@ 10 Mpc       20 Mpc       50 Mpc       100 Mpc
Several studies have assessed the merger rates of MBH binaries as a function of the mass of the merging bodies and redshift [531, Bonetti:2018tpf] employing semi-analytical models for the cosmic evolution of galaxies and MBHs. The obtained predictions can vary substantially depending on the selected black hole seed mass (light versus heavy seeds) and physical prescriptions implemented to describe the physics of the inspiral; the intrinsic number of MBH mergers is typically of order a few tens to a few hundreds per year, with light-seed models predicting a larger number of events dominated by mergers of MBHs with $<10^{4}M_{\odot}$, while the heavy seed models tend to feature fewer events, mainly involving MBHs of $10^{4}-10^{6}M_{\odot}$. The LISA GW detector (which is sensitive to larger masses and higher redshift, if compared with LGWA) will likely have its detection rates (again of tens to hundreds of events per year) dominated by mergers occurring at $z\sim 5$ [Bonetti:2018tpf]; this suggests that the LGWA will be able to detect at least a few (tens) of events, and possibly have its detection rates dominated by lower-$z$ mergers bridging the gap. Possible observations of MBH mergers with LGWA in the mass range $<10^{4}$ M would bring new insights about the rate of low-mass galaxy mergers and the possible collision of massive star clusters (see also Figure 5).
@@
LGWA is going to predominately observe TDEs occurring about relatively low-mass MBHs ($<10^{6}M_{\odot}$), likely inhabiting dwarf galaxies or globular clusters in the local Universe [511]. This poses a series of challenges for the proper modelling of the rates, as the occupation fraction of MBHs in those low-mass systems remains debated [589], as well as their innermost stellar density profile (which remains virtually unresolvable at the scale of a fraction of the MBH influence radius [590], where most of TDEs are expected to be generated [591]). Although earlier works have suggested TDE rates would be larger about low mass MBHs [583, 592], this idea has been recently revised and it seems instead that rates about low-mass MBHs drop significantly over time after an initial burst [591, 593]. Overall, the event rates about low-mass MBHs in the local universe remain under close investigation, and upcoming surveys as LSST/Vera Rubin Observatory as well as ULTRASAT will greatly improve our understanding of this in the next few years, thus allowing to refine the science we expect to obtain from the LGWA.
@@ 3.2.3 GW cosmology
where $\Omega_{m,0}$ and $\Omega_{r,0}$ denote the present-day values of the density fractions for matter and radiation, respectively, while $\rho_{\rm DE}(z)$ is the dark energy density as a function of redshift and $\rho_{0}=3H_{0}^{2}/(8\pi G)$. The crucial aspect in the GW context is then obtaining information on the sources’ redshift. In the best-case scenario, the GW event is followed by a detectable EM counterpart which leads to a precise estimate of $z$, as we will discuss in the following. In case a counterpart is not available, GW detections can still be used for cosmological purposes through statistical methods. As first proposed in [603], it is possible to statistically obtain redshift measurements from potential host galaxies falling within the observed GW’s localization error volumes exploiting galaxy catalogs. Also, a reconstruction of the intrinsic source-frame mass distribution of the observed mergers, in particular in the presence of sharp features, can lead to a statistical determination of the sources’ redshifts, being the observed quantities in the detector-frame redshifted. Several recent studies have successfully tested these methods and their combination, as well as other similar ones, with both real data [613, 617, 627, 614, 616, 615, 618, 633] and simulations [608, 610, 638, 639, 640, 641, 642, 643, 644, 645]. For the former method to be informative, detections with narrow error volumes, both in the angular and distance components, are extremely valuable. Another independent approach to measure the expansion history of the Universe as well as the GW bias parameter is using the cross-correlation technique, which can measure the spatial clustering of GW sources with the galaxies [646, 620, 611, 622, 647, 648, 649, 650, 626, 651]. This method can further explore weak gravitational lensing [621, 652, 653] and provide tests of GR modifications at cosmological scales in a model-independent way [648]. Other methods to employ GW detections for cosmological purposes include, e.g., exploiting multiple images of strongly lensed events [654, 655, 656]; having prior knowledge on the redshift distribution of the observed sources [657, 658]; using tidal distortions of NSs combined with EoS knowledge to extract information on the source-frame masses [607, 659, 660]; or measuring the angular barion acoustic oscillation scale from a high number of well localized events [661].
@@
For GW-only events, the so-called ‘dark sirens’ without an EM counterpart and the host galaxy identification, statistical approaches can be adopted. Despite the low statistics, it has also been possible to get an interesting accuracy on the local expansion rate of the Universe [613, 614, 615, 616, 617, 618], as well as other modified gravity observables [617, 631, 634, 632]. As an example, from the binary black-hole merger GW170814 in combination with the Dark Energy Survey it has been possible to set the constraint $H_{0}=75\begin{subarray}{c}+40\\ -32\end{subarray}~{}{\rm km\,s}^{-1}\,{\rm Mpc}^{-1}$ [614]. Using the sources in the GWTC-3 catalog, the LVK Collaboration obtained a constraint of $H_{0}=50\begin{subarray}{c}+37\\ -30\end{subarray}~{}{\rm km\,s}^{-1}\,{\rm Mpc}^{-1}$ adopting the statistical method relying on the source-frame mass distribution reconstruction and a Power-Law + Peak profile, while a tighter constraint of $H_{0}=67\begin{subarray}{c}+13\\ -12\end{subarray}~{}{\rm km\,s}^{-1}\,{\rm Mpc}^{-1}$ has been obtained from the same sources with the statistical host identification technique [618]. [^12] Notice that this value strongly depends on the assumptions for the BBH mass distribution. Combining with the posterior obtained for GW170817 and its associated counterpart, these constraints shrink to $H_{0}=68\begin{subarray}{c}+12\\ -8\end{subarray}~{}{\rm km\,s}^{-1}\,{\rm Mpc}^{-1}$ and $H_{0}=68\begin{subarray}{c}+8\\ -6\end{subarray}~{}{\rm km\,s}^{-1}\,{\rm Mpc}^{-1}$, respectively. A more recent result obtained in [644] from the combination of these two techniques applied to the GWTC-3 events is $H_{0}=69\begin{subarray}{c}+12\\ -7\end{subarray}~{}{\rm km\,s}^{-1}\,{\rm Mpc}^{-1}$.
@@
In particular, in the bright siren case, we expect EM counterparts from CBCs in the Hz-kHz band to be produced by mergers involving at least one deformable object, such as a neutron star, thus coalescences at $\gtrsim{\cal O}({\rm kHz})$ frequencies, falling outside the LGWA band. However, LGWA is able to detect the early inspiral of such systems and can be crucial to localize the source in the sky with good accuracy, hence improving the chances of counterpart identification (see Sec. 2.2.4). Even though the merit of joint GW+EM detections for cosmology is particularly high, systems with an associated counterpart are expected to be rare, with ${\cal O}(1)$ forecasted BNS multimessenger detections per year with LGWA (see Sec. 2.2.1). It should be noted that, even though the rate of multimessenger BNS events with LGWA is expected to be considerably lower than what will be achievable with 3G ground-based detectors, with tens to hundreds of expected multimessenger detections per year (see e.g. [665, 66]), the improved chances of counterpart identification thanks to the good LGWA localization can be valuable and will raise the number of bright siren events. Moreover, multiband observations of this kind of systems with LGWA and 3G detectors could be a powerful tool to study deviations from GR at the level of propagation of GWs through the Universe [87], with the possible counterpart identification adding even more information as compared to the BBH case.
@@
In the dark siren case instead, at least for some of the coalescences falling in the band of terrestrial detectors, in particular at low redshifts, LGWA can play a crucial role, significantly shrinking the sky localization area thanks to the long time spent by the sources in the dHz detector band, and thus the huge effective baseline (see Sec. 2.2.4). Combined with the accuracy in particular third generation terrestrial interferometers can have in the reconstruction of the luminosity distance, such detections can be a game changer for dark-siren GW cosmology, reducing the number of potential host galaxies in such small error volumes to only a few, and thus acting effectively almost as bright sirens. Moreover, the detections of MBH and IMBH mergers that occur in the LGWA band could be used for both the aforementioned methodologies, increasing the statistics and thus the constraining power, with the statistical uncertainty scaling as $\propto 1/\sqrt{N_{\rm events}}$. Another interesting opportunity offered by these massive sources is the possible detection of EM radiation preceding or accompanying the gravitational signal from the environment surrounding the compact objects, making them bright siren candidates highly valuable for cosmology (see Sec. 3.3.4). Finally, as shown in Fig. 7, LGWA will be able to detect coalescences of DWD and NSWD binaries up to $\sim\!100~{}{\rm Mpc}$. Thanks to the long time this kind of sources spend inspiralling in the detector’s band, they can potentially be localized with very high accuracy, and provide a powerful tool to estimate the $H_{0}$ value in the local Universe, both as bright (see Sec. 3.3.1 for a discussion regarding the EM transients these systems can give rise to) and dark sirens in combination with galaxy catalogs if counterpart detections are not available. The estimation of all the other cosmological parameters impacting the $d_{L}-z$ relation at higher redshifts coming from the other classes of sources discussed above would greatly benefit from this low redshift measurement.
@@
Another observable and yet elusive source for current detectors is a SGWB of cosmological origin (CGWB). This refers to tensor modes possibly produced by various physical processes happening in the early universe, including, e.g., phase transitions, inflation, cosmic strings or primordial black holes (see e.g. [666, 667, 668] for reviews). It is, in a sense, the GW equivalent of the Cosmic Microwave Background EM radiation, whose discovery in 1965 has been one of the most relevant in cosmology. The detection of a CGWB would also bring outstanding scientific value, providing direct information about our Universe in its very first instants of evolution that would not be accessible by other means. [^13] As studied for example in [669], an astrophysical GW background generated by stellar origin CBCs might dominate the SGWB also in the decihertz band. Sophisticatd methods will have to be applied to reduce its contribution for CGWB observations [670, 671, 672]. The characteristic frequency $f_{0}$ of a CGWB at present time is determined by the Hubble rate at the time of generation of the background, redshifted to the present time [667]
@@ 3.3 Multi-messenger observations
The LGWA science case is particularly rich in astrophysical and matter phenomena. As a consequence, many of the GW sources are expected to emit detectable EM signals as well. Most of these events are specific to the deci-Hertz band. Some of these are low-probability events that may be extremely rewarding to detect. In this section, we outline the multi-messenger context of LGWA signals. The prospects of EM observations of binaries containing WDs is described in Sec. 3.3.1. The detection of SNe together with associated GW signals would be a major breakthrough (Sec. 3.3.2). Rare signals observed in EM and GW might come from TDEs of WDs near the horizon of BHs (Sec. 3.3.3). Another multi-messenger study concerns MBH binaries in AGNs, where X-ray observations can complement the GW detection (Sec. 3.3.4). Finally, a new class of EM signals called QPEs are best explained by X-ray emissions from gas in E/IMRI systems (Sec. 3.3.5). We refer to section 2.2.2 for an estimate of LGWA’s sky-localization capabilities, which plays a crucial role for most multi-messenger studies.
@@ 3.3.1 Compact binaries with white dwarfs and neutron stars
Mergers involving double WD binaries and NSWD binaries can lead to explosive or transient events and the formation of exotic objects that typically do not emerge from the evolutionary path of single stars. Many studies have delved into the mergers of double WD binaries, mostly in the context of potential SN-Ia progenitors, as reviewed by various authors such as [824, 825, 826, 275, 827]. These binaries are created through common envelope evolution in binary systems, and in the closest double WD binaries, the emission of GWs causes their orbits to shrink and initiates mass transfer between the WDs within a Hubble timescale. The critical factor influencing the final result is whether mass transfer remains stable or diverges, potentially leading to the tidal disruption of the donor star and binary merger [828]. Mass transfer’s stability depends on the response of the WDs to the process, their masses, and various angular momentum transport mechanisms, such as torques due to accretion discs or tidal bulges [829, 830]. Subsequent evolution after WD mergers can yield diverse outcomes, including He-rich hot subdwarfs, R Coronae Borealis stars, massive carbon/oxygen or oxygen/neon WDs, SN-Ia, accretion-induced collapse to a neutron star or even a black hole. A chart summarizing these potential outcomes of WD mergers can be found in Fig. 3 of [831]. The case of SN Ia holds broader historical significance and particular scientific interest for LGWA, and is therefore discussed in detail in Sec. 3.3.2.
@@
Mergers involving NSWD and BHWD binaries are also expected to fall within the LGWA sensitivity band and generate electromagnetic transients. However, simulating theoretical predictions about their observational signatures to estimate them numerically presents more challenges. The main challenge lies in dealing with a wide range of scales and complex physical processes inherent in this problem. This is in contrast to mergers involving objects of similar size, such as DWD or NSNS/BH, which are more manageable and have more well-developed observational predictions, as demonstrated in studies like [832, 833, 505]. The merger events involving NS or BH with a WD have been extensively studied through various hydrodynamical and nuclear-hydrodynamical simulations. Some notable references in this research include [271, 272, 834, 277, 279]. It is generally anticipated that all NSWD mergers result in unstable mass transfer [274]. In such case, the WD is tidally disrupted on dynamical timescales, forming an extended debris disk around the NS. The evolution of this disk is primarily governed by viscosity, but nuclear burning also plays a significant role. Together, disk viscosity and nuclear burning drive outflows throughout the disk [271]. Nuclear burning within the disk progresses steadily, potentially with a weak detonation, and it results in the production of small amounts of 56Ni, typically at most 10-2 M [834]. This limited amount of 56Ni is expected to lead to optical transients, but they are significantly fainter compared to typical type Ia supernovae, which generate over an order of magnitude more 56Ni. The transients resulting from NSWD mergers could potentially represent an entirely distinct class of supernovae. These events might be observable primarily in nearby galaxies using large telescopes or, perhaps, with next-generation survey instruments like Rubin-LSST.
@@
Binaries with a WD component are not loud emitters of gravitational waves, due to their low masses and long orbital period. Most such binaries will be detected in their inspiral phase integrating the quasi-monotonic signal over years, similarly to what is done for continuous wave searches by current GW detectors. A few of these detected binaries will merge during the LGWA mission lifetime. The presence of eccentricity – expected for NSWD binaries if the NS forms after the WD [275] – gives rise to higher-frequency harmonics of the emission [835, 703], which boost overall emitted power as well as placing it closer to the most sensitive band for LGWA. As a summary of what is discussed in Sec. 2.2.1, deci-Hz GW sources in our own Milky Way are very likely to be observed regardless of their orientation, while sources outside our galaxy might be seen mostly if they have favorable orientations and masses.
Multi-messenger detections of WD mergers within our galaxy should not be expected, even though the Milky Way hosts an estimated $\mathcal{O}(10^{8})$ DWD and $\mathcal{O}(10^{7})$ NSWD binaries [481]. This is because only a small fraction exists as short-period binaries.angesHence, LGWA’s capability to see binaries with WDs out to a few 10 Mpc (see section 2.2.1) is crucial for the success of a EM-GW multi-messenger campaign. For reference, the detection rates of DWD merger derived from SN Ia observations are estimated in section 3.2.2. More details of a DWD and SN Ia association and a possible coincident detection can be found in section 3.3.2.
Nevertheless, pinpointing even a limited number of ultra-short period DWD/NSWD binaries can yield significant astrophysical insight. Optical follow-ups could be essential in resolving parameter ambiguities that arise from relying solely on GW signals. These parameters include system mass, distance, and orbital period, especially if light-curve modulations are discerned. Collectively, such observations can elucidate various domains, from white dwarf structures and binary stellar evolution to accretion physics, general relativity, and the overarching structure of our galaxy. The association of a gravitational event with an electromagnetic counterpart, such as SN-Ia explosion, would be a groundbreaking discovery, much like for the first multi-messenger observation of binary NS merger GW170817 [836].
Concerning NSWD mergers, in the event the NS does not collapse into a BH post-merger, Morán-Fraile et al [279] estimate jet velocities similar to those seen in galactic microquasars ($\gtrsim 0.9c$), leading to an event spectrum that initiates with a soft X-ray or UV flash and transitions into lower frequencies, resembling fast blue optical transients (FBOTs, [837]), potentially associated with NSWD mergers. The merger could generate a prolonged, bright radio signal, with the transient’s luminosity potentially exceeding typical kilonovae due to the ejecta’s mass and velocity, rather than r-process yields, which are obscured in kilonovae by lanthanides. If the NS transitions into a BH, this could result in ultrarelativistic jet speeds, leading to long or ultra-long GRB-like transients, marking a significant observational shift